# D2ST-Adapter: Disentangled-and-Deformable Spatio-Temporal Adapter for Few-shot Action Recognition

### v3

@@ 4.3 Ablation study
RFPL [47] 47.0 61.0 74.6 86.8
OTAM [2] MoLo [43] 55.0 69.6 73.8 85.1
@@
HyRSM [44] 54.3 69.0 73.7 86.1
Bi-MHM [44] MoLo [43] 56.6 70.6 74.0 85.6
D2ST-Adapter 57.0 73.6 75.0 87.0
Effect of configuration of the sampling kernel of aDSTA. To investigate the effect of configuring specialized sampling kernels for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway respectively, we additionally configure an aDSTA-Uniform module, which performs sampling with uniform density in both the spatial and temporal domains. Then we compare the performance of our D2ST-Adapter using two different settings, including 1) ‘aDSTA-Uniform’, in which aDSTA-Uniform is used to model both the spatial and temporal pathways and 2) ‘aDSTA-S & aDSTA-T’, in which D2ST-Adapter uses aDSTA-S and aDSTA-T to model the spatial and temporal pathways respectively. As shown in Table 4, the distinct performance gap between these two settings reveals the advantage of configuring specialized sampling kernels for the spatial and temporal pathways.
@@
OTAM [2] 39.4 45.0 46.6 47.4 49.0 40.6 45.2 48.0 48.9 49.2
TRX [28] 29.9 38.2 44.0 48.2 50.3 32.4 41.6 47.7 52.0 53.5
HyRSM [44] 46.6 54.7 58.7 60.7 61.1 50.0 57.5 61.9 63.3 64.8
MoLo [43] 50.0 57.2 61.6 63.6 64.6 54.1 61.1 65.9 67.3 67.8
ST-Adapter [25] 47.3 54.3 58.1 61.3 62.2 48.9 55.6 59.4 62.5 64.1
@@
Instantiation on ResNet-18 and ResNet-34. To investigate the performance of our D2ST-Adapter on smaller backbones, we instantiate our model with ResNet-18 and ResNet-34 pre-trained on ImageNet, and conduct experiments to compare our model with other methods under the same experimental settings. Table 6 show that our D2ST-Adapter achieves the best performance using both backbones in all settings, which demonstrates the robustness of our method. While DST-Adapter, the convolutional version of our model, also performs well, it is still inferior to D2ST-Adapter, which manifests the effectiveness of the proposed anisotropic Deformable Spatio-Temporal Attention (aDSTA).
Figure 5: Visualization of shifted reference points in both pathways for two video samples. Circle size indicates the importance for each point calculated by aggregating the attention from all queries.
4.4 Visualization
We visualize the shifted reference points learned by our D2ST-Adapter to investigate the effectiveness of the proposed aDSTA qualitatively. Following Deformable Attention [48], for each shifted reference point serving as key and value, we accumulate the attention weights assigned by all queries as its relative importance. Then we visualize top-100 most important shifted reference points in each pathway for a video sample, as shown in  Figure 5. Note that we sample 8 frames for all videos, if a reference point is shifted to a temporal position between two sampled frames, we visualize it on the nearest frame approximately.
The results show that the shifted reference points learned by our model can always focus on the salient targets that are critical for action recognition in both the spatial and temporal pathways. Another interesting observation is that the shifted points in the spatial pathway are distributed densely in some specific frames to capture the spatial appearance features for the targets, while the shifted points in the temporal pathway are distributed uniformly among all frames to capture the temporal dynamic features. This is consistent with the design of specialized tailored aDSTA modules for different pathways by configuring the sampling kernels accordingly.
5 Conclusion
We present D2ST-Adapter, which is a novel adapter tuning method for few-shot action recognition. It is designed in a dual-pathway architecture, which allows for disentangled encoding of spatial and temporal features. Moreover, we design the anisotropic Deformable Spatio-Temporal Attention (aDSTA) as the essential component of D2ST-Adapter, which enables it to encode features in a global view while maintaining lightweight design. Extensive experiments demonstrate that the superiority of our model over state-of-the-art methods, especially in challenging scenarios involving complex temporal features.
Supplementary Material
Appendix 0.A Tuning of the Sampling Kernels of aDSTA
The sampling kernels of aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter can be tuned on a held-out validation set. Generally, aDSTA-S should sample denser reference points in the spatial domain while aDSTA-T samples more points in the temporal domain. For the instantiation of our model with ResNet, we configure the sampling kernel of aDSTA in each convolutional stage individually since the feature maps in different convolutional stages may have different size. In contrast, we only need to tune one configuration for sampling kernel when using ViT as the backbone since the feature map always has fixed size in different stages. Table 7 shows the tuned configurations of the sampling kernels of aDSTA-S and aDSTA-T with different backbones. Besides, we also provide the configurations of sampling kernel for aDSTA-Uniform module which is constructed to validate the effect of configuring the sampling kernel of aDSTA in Table 4.
Table 7: Tuned sampling kernels (in the form of $(T,H,W)$) for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter. Besides, the sampling kernel for aDSTA-Uniform used in Table 4 for ablation study is also provided.
Backbone Feature Map Sampling Kernel Sampling Kernel Sampling Kernel
@@ Appendix 0.A Tuning of the Sampling Kernels of aDSTA
ResNet-50 (8, 56, 56) (2, 8, 8) (8, 4, 4) (4, 4, 4)
ResNet-50 (8, 28, 28) (2, 4, 4) (8, 2, 2) (4, 4, 4)
ResNet-50 (8, 14, 14) (2, 4, 4) (8, 2, 2) (4, 4, 4)
ResNet-50 (8, 7, 7) (2, 2, 2) (8, 1, 1) (4, 4, 4)
CLIP-ViT-B (8, 14, 14) (2, 4, 4) (8, 2, 2) (4, 4, 4)
Appendix 0.B Effect of the Inserted Position of D2ST-Adapter
Theoretically, our D2ST-Adapter can be inserted into any position of the backbone flexibly. To investigate the effect of the inserted position of D2ST-Adapter on the performance of the model, we conduct experiments with four different ways of inserting D2ST-Adapters into the pre-trained CLIP-ViT-B backbone (comprising 12 learning stages) on SSv2-Small dataset: a) early-insertion, which inserts the D2ST-Adapter into each of first 6 stages (close to the input), b) late-insertion that inserts the D2ST-Adapter into each of last 6 stages (close to the output), c) skip-insertion, which inserts the adapter into the backbone every two stages and d) full-insertion that inserts the adapter into each learning stage.
As shown in Table 8, late-insertion, namely inserting the proposed D2ST-Adapters into the last 6 stages, yields better performance than early-insertion, which implies that adapter tuning is more effective for task adaptation in deeper layers than in the shallower layers. It is reasonable since deeper layers generally capture the high-level semantic features, which are more relevant to task adaptation. Besides, full-insertion achieves the best performance at the expense of slightly more tunable parameters and memory usage.
Insertion Position Tunable Parameters (%) Memory Usage 1-shot 5-shot
Early-insertion 4.1% 15.5GB 48.4 64.6
Late-insertion 4.1% 15.2GB 54.2 68.5
Skip-insertion 4.1% 15.2GB 53.3 67.9
Full-insertion 7.9% 16.7GB 55.0 69.3
Ratio Tunable Parameters (%) Memory Usage 1-shot 5-shot
0.0625 1.4% 15.2GB 53.4 68.1
0.125 3.2% 15.7GB 54.2 68.6
0.25 7.9% 16.7GB 55.0 69.3
0.5 20.2% 18.7GB 54.7 69.4
Appendix 0.C Effect of the Bottleneck Ratio of D2ST-Adapter
@@
Figure 6: Given a selected query within the salient object, top-50 relevant shifted reference points in terms of attention weights in both pathways are visualized for three video samples. The red stars denote the selected query tokens while the circles denote relevant shifted reference points.
Appendix 0.D More Visualizations
To provide more insight into the anisotropic Deformable Spatio-Temporal Attention (aDSTA), we provide more visualization results in Figure 6. To be specific, we manually select a point within the salient object as the query token and visualize top-50 relevant shifted reference points to the query in terms of attention weights. The results show that our model is able to capture the salient object through the shifted reference points in both spatial and temporal domains.

### v4

@@ 4.3 Ablation study
RFPL 47.0 61.0 74.6 86.8
OTAM MoLo 55.0 69.6 73.8 85.1
@@
HyRSM 54.3 69.0 73.7 86.1
Bi-MHM MoLo 56.6 70.6 74.0 85.6
D2ST-Adapter 57.0 73.6 75.8 87.7
Visualization. We visualize the shifted reference points learned by our D2ST-Adapter to investigate the proposed aDSTA qualitatively. Following Deformable Attention [40], for each shifted reference point serving as key and value, we accumulate the attention weights assigned by all queries as its relative importance. Then we visualize top-100 most important shifted reference points in each pathway, with circle size indicating their importance, as shown in the first sample of Figure 5. We also provide visualization of the most relevant reference points w.r.t a certain salient region (token) in the second sample of Figure 5. To be specific, we manually select a salient region (indicated by a red star marker) as the query token and visualize top-50 relevant shifted reference points to the query in terms of attention weights.
The results show that the shifted reference points learned by our model can always focus on the salient targets that are critical for action recognition in both the spatial and temporal pathways. Another observation is that the shifted points in the spatial pathway are distributed densely in some specific frames to capture the spatial appearance features for the targets, while the shifted points in the temporal pathway are distributed uniformly among all frames to capture the temporal dynamics. This is consistent with the anisotropic sampling of aDSTA for different pathways.
5 Concluding Remarks
We present D2ST-Adapter, which is a novel adapter tuning method for few-shot action recognition. It is designed in a dual-pathway architecture, which allows for disentangled encoding of spatial and temporal features. Moreover, we design the aDSTA module, which enables D2ST-Adapter to encode features in a global view while maintaining lightweight design. Extensive experiments demonstrate the superiority of our model over state-of-the-art methods, especially in challenging scenarios involving complex temporal features.
Supplementary Material
Appendix A Effect of the Sampling Densities of D2ST-Adapter
We compare the efficiency and performance of different sampling densities in 1-shot setting in Table 7. It demonstrates that sampling denser reference points yields marginal performance gain at the expense of increased memory usage. Our model is set to $<$2, 4, 4$>$ &\&& $<$8, 2, 2$>$ to strike a good balance between efficiency and performance.
Appendix B Effect of the Number of Input Frames
To analyze the impact of the number of input frames, we compare the SSv2-Full 1-shot performance when using different number of frames as input under two sampling densities settings. As shown in Table 8, more input frames and denser sampling both yield better performance while incurring more computational cost.
Appendix C Instantiation on More Backbones
To investigate the performance of our D2ST-Adapter on smaller backbones, we instantiate our model with ResNet-18 and ResNet-34 pre-trained on ImageNet, and conduct experiments to compare our model with other methods under the same experimental settings. Table 11 shows that our D2ST-Adapter achieves the best performance using both backbones in all settings, which demonstrates the robustness of our method across different backbones. While DST-Adapter, the convolutional version of our model, also performs well, it is still inferior to D2ST-Adapter, which manifests the effectiveness of the proposed anisotropic Deformable Spatio-Temporal Attention (aDSTA).
For larger backbones, we conduct experiments on CLIP-ViT-L/14, which has a total of 303.2 million parameters, significantly larger than CLIP-ViT-B/16, which has a total of 85.8 million parameters. The results in Table 12 show that our model can still consistently outperform ST-Adapter and Full Fine-tuning on the larger and stronger backbone.
Table 7: Effect of the sampling densities of D2ST-Adapter in CLIP-ViT-B.
Sampling Densities Memory Usage SSv2-Full Kinetics
(Spatial & Temporal)
$<$1, 4, 4$>$ &\&& $<$4, 2, 2$>$ 17.0 GB 65.9 89.0
$<$2, 4, 4$>$ &\&& $<$8, 2, 2$>$ 17.3 GB 66.7 89.3
$<$2, 6, 6$>$ &\&& $<$8, 3, 3$>$ 18.0 GB 66.9 89.2
$<$2, 8, 8$>$ &\&& $<$8, 4, 4$>$ 18.8 GB 66.9 89.6
$<$4, 8, 8$>$ &\&& $<$16, 4, 4$>$ 20.3 GB 67.2 89.3
Table 8: Effect of the number of input frames in SSv2-Full 1-shot setting with CLIP-ViT-B backbone.
Sampling Densities Sampling Frames
(Spatial & Temporal) 4 8 16
$<$1, 4, 4$>$ &\&& $<$4, 2, 2$>$ 63.1 65.9 69.0
$<$2, 4, 4$>$ &\&& $<$8, 2, 2$>$ 63.6 66.7 69.9
Insertion Position Tunable Memory Usage 1-shot 5-shot
Params (%)
Early-insertion 4.1% 15.7 GB 48.4 64.6
Late-insertion 4.1% 15.3 GB 54.2 68.5
Skip-insertion 4.1% 15.4 GB 53.3 67.9
Full-insertion 7.9% 17.3 GB 55.0 69.3
Ratio Tunable Memory Usage 1-shot 5-shot
Params (%)
0.0625 1.4% 15.5 GB 53.4 68.1
0.125 3.2% 16.2 GB 54.2 68.6
0.25 7.9% 17.3 GB 55.0 69.3
0.5 20.2% 19.7 GB 54.7 69.4
@@ Appendix C Instantiation on More Backbones
OTAM 39.4 45.0 46.6 47.4 49.0 40.6 45.2 48.0 48.9 49.2
TRX 29.9 38.2 44.0 48.2 50.3 32.4 41.6 47.7 52.0 53.5
HyRSM 46.6 54.7 58.7 60.7 61.1 50.0 57.5 61.9 63.3 64.8
MoLo 50.0 57.2 61.6 63.6 64.6 54.1 61.1 65.9 67.3 67.8
ST-Adapter 47.3 54.3 58.1 61.3 62.2 48.9 55.6 59.4 62.5 64.1
@@
Table 12: Performance of different methods using a larger backbone (CLIP-ViT-L/14). The number of backbone parameters (Params) is also provided as reference.
Method Backbone (Params) SSv2-Full Kinetics
1-shot 5-shot 1-shot 5-shot
Full Fine-tuning CLIP-ViT-L/14 (303.2 M) 55.5 73.0 91.7 96.8
ST-Adapter CLIP-ViT-L/14 (303.2 M) 67.0 83.8 92.2 96.9
D2ST-Adapter CLIP-ViT-L/14 (303.2 M) 69.6 85.6 92.8 97.2
Table 13: Hyper-parameter settings of our method with ResNet-50 backbone.
Backbone Hyper-parameter SSv2-Full SSv2-Small Kinetics HMDB51 UCF101
ResNet-50 base learning rate 2e-3 2e-3 2e-3 2e-3 2e-3
warmup start learning rate 4e-4 4e-4 4e-4 4e-4 4e-4
training episodes 120000 40000 10000 8000 8000
warmup episodes 12000 4000 1000 800 800
weight decay 5e-4 5e-4 5e-4 5e-4 5e-4
batch size 8 8 8 8 8
num. adapters per block 1 1 1 1 1
num. input frames 8 8 8 8 8
bottleneck ratio 0.25 0.25 0.25 0.25 0.25
training crop size 224 224 224 224 224
CLIP-ViT-B base learning rate 1e-3 1e-3 5e-4 1e-3 5e-4
warmup start learning rate 2e-4 2e-4 1e-4 2e-4 1e-4
training episodes 120000 40000 3000 3000 2400
warmup episodes 12000 4000 300 300 240
weight decay 5e-4 5e-4 5e-4 5e-4 5e-4
batch size 8 8 8 8 8
num. adapters per block 1 1 1 1 1
num. input frames 8 8 8 8 8
bottleneck ratio 0.25 0.25 0.25 0.25 0.25
training crop size 224 224 224 224 224
Table 14: Tuned sampling densities (in the form of $<$T, H, W$>$) for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter. Besides, the sampling densities for aDSTA-Uniform used in Table 5 for ablation study is also provided.
Backbone Feature Map Sampling Densities Sampling Densities Sampling Densities
@@
ResNet-50 $<$8, 56, 56$>$ $<$2, 8, 8$>$ $<$8, 4, 4$>$ $<$4, 4, 4$>$
ResNet-50 $<$8, 28, 28$>$ $<$2, 4, 4$>$ $<$8, 2, 2$>$ $<$4, 4, 4$>$
ResNet-50 $<$8, 14, 14$>$ $<$2, 4, 4$>$ $<$8, 2, 2$>$ $<$4, 4, 4$>$
ResNet-50 $<$8, 7, 7$>$ $<$2, 2, 2$>$ $<$8, 1, 1$>$ $<$4, 4, 4$>$
CLIP-ViT-B $<$8, 14, 14$>$ $<$2, 4, 4$>$ $<$8, 2, 2$>$ $<$4, 4, 4$>$
Appendix D Effect of the Inserted Position of D2ST-Adapter
Theoretically, our D2ST-Adapter can be inserted into any position of the backbone flexibly. To investigate the effect of the inserted position of D2ST-Adapter on the performance of the model, we conduct experiments with four different ways of inserting D2ST-Adapters into the pre-trained CLIP-ViT-B backbone (comprising 12 learning stages) on SSv2-Small dataset: a) early-insertion, which inserts the D2ST-Adapter into each of first 6 stages (close to the input), b) late-insertion that inserts the D2ST-Adapter into each of last 6 stages (close to the output), c) skip-insertion, which inserts the adapter into the backbone every two stages and d) full-insertion that inserts the adapter into each learning stage. As shown in Table 9, late-insertion, namely inserting the proposed D2ST-Adapters into the last 6 stages, yields better performance than early-insertion, which implies that adapter tuning is more effective for task adaptation in deeper layers than in the shallower layers. It is reasonable since deeper layers generally capture the high-level semantic features, which are more relevant to task adaptation. Besides, full-insertion achieves the best performance at the expense of slightly more tunable parameters and memory usage.
Appendix E Effect of the Bottleneck Ratio of D2ST-Adapter
@@
Appendix F More Visualizations
Consistent with Figure 5 in the paper, we provide more visualization of the shifted reference points in Figures 6 and 7 to evaluate the anisotropic Deformable Spatio-Temporal Attention (aDSTA). Note that only the top-100/50 most important points from the aDSTA modules across all inserted D2ST-Adapters are visualized. The results show that our model is able to capture the salient objects through the shifted reference points in both spatial and temporal domains.
Appendix G More Implementation Details
Matching Metric. In the few-shot action recognition task, the classification of a query sample is based on the similarities between it and each support class prototype. Typically, a matching metric is first used to temporally align the frames or segments within two videos. Then, the overall similarity between the query and each prototype can be calculated by averaging or summing the similarities of all aligned pairs. We adopt three classic matching metrics in our experiments, including OTAM [3], TRX [23], and Bi-MHM [33]. Details of these matching metrics are provided as follows.
OTAM [3] explicitly leverages the temporal ordering information in videos through tight temporal alignment, which ensures that the alignment results for any two frame pairs do not overlap. It extends the Dynamic Time Warping (DTW) algorithm [19] to compute the alignment path of two videos in the temporal dimension, and employs continuous relaxation to make the model differentiable.
TRX [23] aligns segments rather than individual frames, thus effectively matching actions performed at varying speeds and in different locations across two videos. It first exhaustively enumerates all subsequences of two to four frames as potential actions in both videos. Then, for each action in the query video, the cross-attention mechanism is employed to compute the corresponding action prototype in the support video based on feature similarities, serving as its match.
Bi-MHM [33] stands for Bidirectional Mean Hausdorff Metric, which treats the similarity measurement between two videos as a set matching problem. This alignment strategy eliminates the constraints of temporal order and focuses solely on the appearance similarities between frames.
Data Pre-processing. We use the same data pre-processing methods as in HyRSM [33]. A video is uniformly sampled 8 frames as input. During training, each frame is first resized to 256$\times$256 and then randomly cropped to 224$\times$224. Some basic data augmentation methods are adopted, such as color jitter and horizontal flip. Note that for temporal-related datasets, i.e. SSv2-Full and SSv2-Small, we do not use horizontal flip, since the recognition of some classes in these two datasets requires distinguishing between left and right, e.g. “pulling something from left to right”.
Tuned hyper-parameters. For ease of the reproduction, we list the tuned values for all the hyper-parameters in our experiments for each dataset with both backbones in Table 13.
Tuned sampling densities of aDSTA. The sampling densities of aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter can be tuned on a held-out validation set. Generally, aDSTA-S should sample denser reference points in the spatial domain while aDSTA-T samples more points in the temporal domain. For the instantiation of our model with ResNet, we configure the sampling densities of aDSTA in each convolutional stage individually since the feature maps in different convolutional stages may have different size. In contrast, we only need to tune one configuration for sampling densities when using ViT as the backbone since the feature map always has fixed size in different stages. Table 14 shows the tuned configurations of the sampling densities of aDSTA-S and aDSTA-T with different backbones. Besides, we also provide the configurations of sampling densities for aDSTA-Uniform module which is constructed to validate the effect of configuring the sampling densities of aDSTA in Table 5.
Figure 6: Visualization of top-100 important shifted reference points in both pathways for two video samples. Circle size indicates the importance for each point calculated by aggregating the attention weights from all queries.
Figure 7: Given a selected query within the salient object indicated by a red star marker, top-50 relevant shifted reference points in terms of attention weight are visualized.